OpenStudy (anonymous):

Please explain how to get my missing part. I have to write a quadratic equation in standard form with the given solution set of {-7/8,1/3}. So I make x+7/8=0 and x-1/3=0 then to 8x+7=0 and 3x-1=0 from that I get (8x+7)(3x-1)=0 and finally I can com up with 24x^2+?x-7=0 I know after getting it wrong that my ?=13 but I do not understand who they got to the conclusion. Please explain. thank you. i got 24 from 8*3 and I for the -7 from 7*-1.

OpenStudy (jdoe0001):

I'm not too sure I understand what you're asking :| your quadratic is correct from the roots give, that is, \(\bf (8x+7)(3x-1)=0\implies 24x^2+13x-7=0\) so.... what is it nagging you?

OpenStudy (anonymous):

How did they come up with the 13x for b I am just not seeing it.

OpenStudy (jdoe0001):

\(\bf (8x+7)(3x-1)=0\\ \quad \\ (8x+7)(3x-1)\implies \begin{cases} 8x(3x-1)\implies 24x^2-8x\\ \quad \\ \bf 7(3x-1)\implies 21x-7 \end{cases}\\ \quad \\ 24x^2-8x+21x-7\implies 24x^2(-8x+21x)-7\\ \implies 24x^2+13x-7\)

OpenStudy (anonymous):

Explained perfectly, thank you